OpenStudy (anonymous):

The price of a notebook has risen to $3.90 today. Yesterday's price was $3.40 . Find the percentage increase. Round your answer to the nearest tenth of a percent.

OpenStudy (anonymous):

$3.90 -$3.40 =$0.5 ($0.5/$3.90 )=(x/100) ( 0.5*100)=3.4x x =50/3.9 x=12.8205 percent